OpenStudy (anonymous):

Solve for x √3x-5=2 A.no solution B.7/3 C.–1 D.3 Please Help

OpenStudy (paki):

root is over 3x only or over 5 too

OpenStudy (anonymous):

over both

OpenStudy (anonymous):

First square both sides & simplify it for x

OpenStudy (paki):

take square on both sides...

OpenStudy (anonymous):

making it \[3x-5=4\]

OpenStudy (paki):

it will be (√3x-5)^2 = (2)^2 ... it will cancel the root... and you will obtain an equation like... 3x-5=4... now solve ofr x...

OpenStudy (anonymous):

then \[3x-5+5=4+5\]

OpenStudy (paki):

yeah

OpenStudy (anonymous):

\[3x \div3=9\div3\]

OpenStudy (paki):

yeah :-)

OpenStudy (anonymous):

\[x=3\]

OpenStudy (paki):

yeah...

OpenStudy (anonymous):

Thanks so much!

OpenStudy (paki):

good :-)

OpenStudy (paki):

pleasure